HETAV PATEL

Grade 11,

a tube well pumps 2400 kg/mins .if the water is coming out with velocity 3m/s.what is the power of the pump?

a tube well pumps 2400 kg/mins .if the water is coming out with velocity 3m/s.what is the power of the pump?

Grade:9

2 Answers

Ramesh V
14 years ago

power =  force* velocity

force exerted on water in 1 sec = F = m*(v/t)

                                                   =(2400/60)*(3/1) = 120 N

power exerted = F*v=120*3=360 W

--

regards

Ramesh

Harish
6 years ago
Mass of water pumped out per sec = 2400/60 = 40 kg/sPower of pump = kinetic energy of water coming out per sec = mv^2/2 = 180 J Harish kaswan